OpenStudy (mackenzie2013):

If you roll a number cube 96 times, how many times would you expect to roll a prime number?

OpenStudy (valpey):

Which of the sides of the number cube are prime?

OpenStudy (amoodarya):

primes={2,3,5} \[p(\left\{ 2,3,6 \right\})=\frac{3}{6}=\frac{1}{2}\] it means average half of them so in 96 times , half of them ...?

OpenStudy (mackenzie2013):

Thank you!